a) Chaque composant est défectueux (S) ou non (É), indépendamment, avec \(p=0{,}03\). \(n=20\) prélèvements. \(X\sim\mathcal{B}(20\,;\,0{,}03)\).
b)
\[\begin{aligned}P(X=0) &= 0{,}97^{20}\\[.35em]&\approx 0{,}5438.\end{aligned}\]
c)
\[\begin{aligned}P(X\leqslant 1) &= P(X=0)+P(X=1)\\[.35em]&= 0{,}5438+\binom{20}{1}(0{,}03)(0{,}97)^{19}\\[.35em]&\approx 0{,}5438+20\times 0{,}03\times 0{,}5607\\[.35em]&\approx 0{,}5438+0{,}3364\\[.35em]&\approx 0{,}8802.\end{aligned}\]
d)
\[\begin{aligned}P(X\geqslant 2) &= 1-P(X\leqslant 1)\\[.35em]&\approx 1-0{,}8802\\[.35em]&\approx 0{,}1198.\end{aligned}\]
e)
\[P(X\geqslant 3)=1-P(X\leqslant 2).\]
\[\begin{aligned}P(X=2) &= \binom{20}{2}(0{,}03)^2(0{,}97)^{18}\\[.35em]&= 190\times 0{,}0009\times 0{,}5780\\[.35em]&\approx 0{,}0988.\end{aligned}\]
\[\begin{aligned}P(X\leqslant 2) &\approx 0{,}8802+0{,}0988\\[.35em]&\approx 0{,}9790.\end{aligned}\]
\[\begin{aligned}P(X\geqslant 3) &\approx 0{,}021\\[.35em]&\approx 2{,}1\%.\end{aligned}\]